OpenStudy (anonymous):

Triangle HIJ is similar to triangle LMN. Side LM measures x + 1, side MN measures 2x and side NL measures 3x – 2. Side HI measures 12 and side IJ measures 20. Find the value of x.

OpenStudy (nurali):

Since ΔHIJ is similar to ΔLMN LM/HI = MN/IJ = NL/JH (x+1)/12 = 2x/20 Cross Multiply 20(x+1) =12*2x 20x + 20 = 24x 4x = 20 x = 5

OpenStudy (anonymous):

Thank you your very helpful I Need Help with One More Does X=4 I Feel Like I Messed Up Right Here 

OpenStudy (anonymous):

help please